OpenStudy (anonymous):

Zach borrowed $4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved $5,920 to pay back to his grandfather. At what interest rate did Zach borrow the money ? a. 23 % b. 16% c. 32 % d. 11%

jimthompson5910 (jim_thompson5910):

What's the interest here?

OpenStudy (anonymous):

im not sure , that is all the info the question gave me ...

jimthompson5910 (jim_thompson5910):

He borrowed 4 thousand and has to pay back 5920, so the interest is 5920 - 4000 = 1920 dollars

OpenStudy (anonymous):

alrigjht , whats next ?

jimthompson5910 (jim_thompson5910):

Since Amount to pay back = principle + interest

jimthompson5910 (jim_thompson5910):

So because the interest is 1920, we know that I = 1920

OpenStudy (anonymous):

and whats the principal ?

jimthompson5910 (jim_thompson5910):

Plug this along with p = 4000 (the principle) and t = 3 (time in years) into I = Prt

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

to get 1920 = 4000*r*3

jimthompson5910 (jim_thompson5910):

now solve for r

OpenStudy (anonymous):

How do i solve for r ? Sorry , im really bad at math :/

jimthompson5910 (jim_thompson5910):

No worries, just keep practicing and you'll get it 1920 = 4000*r*3 1920 = 4000*3*r 1920 = 12000*r 1920/12000 = r 0.16 = r r = 0.16 So the interest rate is 16%

OpenStudy (anonymous):

Thanks again ! (:

jimthompson5910 (jim_thompson5910):

sure thing